When people smoke, the nicotine they absorb is converted to cotinine, which can be measured. A sample of 40 smokers has a mean cotinine level of 172.5. Assuming that
1
Expert's answer
2020-04-02T13:25:36-0400

When people smoke, the nicotine they absorb is converted to cotinine, which can be measured.

A sample of 40 smokers has a mean cotinine level of 172.5. Assuming that "\sigma" is known to be 119.5, find a 90% confidence interval estimate of the mean cotinine level of all smokers.


Solution

We need to construct the "90\%" confidence interval for the population proportion. We have been provided with the following information: "\bar{X}=172.5,\sigma=119.5, n=40, \alpha=0.1."

The critical value for "\alpha=0.1"  is "z_c=z_{1-\alpha/2}=1.645." The corresponding confidence interval is computed as shown below:


"(CI)=(\bar{X}-z_c\times{\sigma\over \sqrt{n}},\bar{X}+z_c\times{\sigma\over \sqrt{n}})="

"=(172.5-1.645\times{119.5\over \sqrt{40}},172.5+1.645\times{119.5\over \sqrt{40}})="

"=(141.418, 203.582)"

Therefore, based on the data provided, the "90\%" confidence interval for the population mean is "(141.418, 203.582)," which indicates that we are 90% confident that the true population mean "\mu"

is contained by the interval "(141.418, 203.582)."